OpenStudy (anonymous):

Graph the circle (x+2)^2+(y+1)^2=32.Label the center and at least 4 points on the circle. Please help! I think the center is (-2,-1) but I don't know how to get the points

OpenStudy (anonymous):

@mathstudent55 help?

OpenStudy (mandre):

You read the center of the circle from the equation: (x - a)^2 + (y - b)^2 = r^2 (a,b) will be the center where (a,b) are the (x,y) coordinates. Remember the minuses in the above formula. In your case it will be (x-(-2))^2 + (x-(-1)^2 = 32 Take the Square root of 32 for your radius and there you have your circle. Substitute x=0 into your formula to work out the y intercepts. Remember there will be 2 (if any, the circle can be away from the axes depending on the center). Substitute y=0 for your x-intercepts. You were actually almost there. You had the center. You just needed the radius.

OpenStudy (mandre):

You don't need the intercepts but it helps to check if your graph is in the right place.